I am three years older than my sister;our ages add up to 21
Answer by Cromlix(4381)

Hi there,
I am three years older than my sister;
our ages add up to 21
I = S + 3.......(1)
I + S = 21......(2)
Substitute (S + 3) for I
in Equation (2)
I + S = 21
(S + 3) + S = 21
S + 3 + S = 21
Collect like terms:
S + S = 21 - 3
2S = 18
S = 9
You are 12 years old
Your sister is 9 years old
Hope this helps :-)